OpenStudy (anonymous):

radical 24 divided by radical 6

OpenStudy (anonymous):

\[\frac{ \sqrt{24} }{ \sqrt{6} } \implies \sqrt{\frac{ 24 }{ 6 }} \implies \sqrt{4} = ?\]